150ml of a gas at 27C is heated to 77C at constant pressure.find the new volume of the gas.
1
Expert's answer
2012-10-31T09:35:09-0400

(V1/T1)= (V2/T2)
V2 = (T2 * V1)/T1
V2 =(350 K0*150ml)/300 = 175 ml